OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the point (7, 8) and has a slope of −1.2. y = x +

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Okay, for this one we plug it into point-slope form and simplify it. \(y - y_1 = m(x - x_1)\) Can you plug in the slope and the point into here?

OpenStudy (anonymous):

y-8=-1.2(x-7) @iGreen

OpenStudy (igreen):

Yes, now we distribute -1.2 into the parenthesis: y - 8 = -1.2(x - 7) y - 8 = -1.2x + 8.4 Add 8 to both sides, what's 8.4 + 8?

OpenStudy (anonymous):

16.4

OpenStudy (igreen):

Yes, so we have: y = -1.2x + 16.4

OpenStudy (igreen):

That's your final answer.

OpenStudy (anonymous):

thank you

OpenStudy (igreen):

Np